Find the minimum distance between the curves y^2 = x-1 and x^2 = y-1                                   please give detail  solution


Grade:12

1 Answers

Ramesh V
14 years ago

Interesting problem.

I found the distance is from point ( 5/4 , 1/2 ) to point ( 1/2, 5/4 )


Note the symmetry of the curve.


So the minimum distance is (3/4) * sqrt (2 ) which is approximately 1.06


I will come back with more details.


My general approach is this:


let the points be (x1,y1) and (x2,y2 )


Find the distance square


d^2 = (x1 -x2) ^2 + (y1 -y2)^2


eliminate (y1, y2 ) from d^2


set ( d (d^2) / dx1 =0 , d (d^2) /dx2 =0 )


solve for (x1,x2)

--

regards

Ran\mesh